The intensity of sound gets reduced by 20% when passed throgh a slab. The reduction in intensity when passed through two consecutive slabs is 
a)40%
b)36%
c)50%
d)30%

Grade:Select Grade

1 Answers

Aziz Alam IIT Roorkee
9 years ago
Initial intensity= I0
Instensity after passing through the 1stslab= 0.8*I0
Intensity after passing through the 2ndslab= 0.8*(0.8*I0)=0.64*I0
SO it can be seen that the intensiity took a hit by 36%